Logiweb codex of factorial

Up Help factorial, ∗!,

factorial

[factorial →^{pyk} “factorial”]

∗!

[x! →^{val} If(x ≈ 0, 1, x · x − 1!)]

[x! →^{tex} “#1.!”]

The pyk compiler, version 0.grue.20050502 by Klaus Grue GRD-2005-06-29.UTC:01:03:50.943925 = MJD-53550.TAI:01:04:22.943925 = LGT-4626723862943925e-6